OpenStudy (anonymous):

Solve by completing the square. x2=3−3x

sam (.sam.):

x2=3−3x x2+3x-3=0 follow my steps, divide 2 for mid, x2+(3x/2)-3=0 square the middle term then bring it outside, add a minus to the one you squared (x+3/2)^2-9/4-3=0 (x+3/2)^2-21/4=0

sam (.sam.):

\[\left( \begin{array}{c} \left( \begin{array}{c} \ (\text{Add } \frac{21}{4} \text{ to both sides}) \\ \left(x+\frac{3}{2}\right)^2=\frac{21}{4} \\ \end{array} \right) \\ \left( \begin{array}{c} \text{Take the square root of both sides} \\ \left(x+\frac{3}{2}=\frac{\sqrt{21}}{2}\right) \text{ or } \left(x+\frac{3}{2}=-\frac{\sqrt{21}}{2}\right) \\ \end{array} \right) \\ \left( \begin{array}{c} (\text{Subtract } \frac{3}{2} \text{ from both sides}) \\ \left( \begin{array}{c} x=\frac{\sqrt{21}}{2}-\frac{3}{2} \\ \end{array} \right) \text{ or } \left(x=-\frac{\sqrt{21}}{2}-\frac{3}{2}\right) \\ \end{array} \right) \\ \end{array} \right)\]

OpenStudy (anonymous):

ohhh ok thank u!! :)